# Belobog: Move Language Fuzzing Framework For Real-World Smart Contracts

### v1

@@ 8. Conclustion
In this paper, we present Belobog, the first effective fuzzing framework for Move smart contracts. We identified three key challenges in generating valid transactions for Move smart contracts, i.e., type safety of objects, type parameters of functions, and type abilities of objects. In order to tackle these challenges, we propose a type graph guided fuzzing approach, where we use a type graph to model all the type constraints of the target smart contract. Upon fuzzing starts, Belobog will query the type graph to generate or mutate transactions that satisfy all the type requirements. To further tackle the strict checks that are common in smart contracts, we design and implement a conclic executor in Belobog.

### v2

@@ 8. Conclusion
In this paper, we present Belobog, the first effective fuzzing framework for Move smart contracts. We identified three key challenges in generating valid transactions for Move smart contracts, i.e., type safety of objects, type parameters of functions, and type abilities of objects. To address these challenges, we propose a type graph-guided fuzzing approach, in which we use a type graph to model all type constraints of the target smart contract. When fuzzing starts, Belobog will query the type graph to generate or mutate transactions that satisfy all the type requirements. To further tackle the strict checks that are common in smart contracts, we design and implement a concolic executor in Belobog. Our extensive evaluation on two real-world datasets demonstrates the capability of Belobog.
9. Data Availability
Due to the page limit, we provide a bundle of supplementary materials.
  • A full copy of Belobog source code.
  • The implementation details of our concolic executor in Section 3.5.
  • Additional case studies of the evaluation of Belobog, including reproducing Nemo, a \$2.59 million incident.
  • Anonymized contents of DS1 and DS2.
  • Detailed evaluation breakdowns for DS1 and DS2.
Upon acceptance, we will fully release Belobog.